OpenStudy (johnnyrhitter):

6+9[10-2(2+3)] What are the properties used in each step?

OpenStudy (jmccollough):

it equals 6

OpenStudy (jmccollough):

= 6+9[10-2(2+3)] = 6+9*(10-2*(2+3)) = 6+9*(10-2*5) = 6+9*0 = 6

OpenStudy (mathmale):

@jmccollough : Next time, would you please guide the other person through the solution of his or her posted problem, instead of solving the problem yourself? Thank you.

OpenStudy (mathmale):

There's more to this problem than just the answer. One must explain which property is used during each step of the problem solving process. In 6+9[10-2(2+3)], the work inside parentheses, (2+3), MUST be done first, followed by the indicated multiplication, followed by ... what? Please review "order of operations" rules.